UT 11 December 2019 - Engineering Night (primary goal: ESM)

OSA: D.G.Huerta

ISp: C. Bolyard

Mtn Manager: B. Bingham

ISA: B. Rothberg (Tucson)

AO: G. Taylor (Tucson)
Software: I. Sola (Tucson - on-call)

NOTES: AGW-4 "wonky" - suggestion is to not use for transform, use at-our-own-risk, if error occurs stop immediately


UT 23:55 (-1 day) - starting up LUCI GUIs

-running init all on both

UT 00:02 - taking test darks -

luciN.20191211.0001-0002 - 2.51sec Darks

OSCO

UT 00:38 - SX OSCO checkout - preset failed, we were authorized as binocular, preset is L1

UT 00:40 - doing pointing check first

UT 00:54 - took some time to collimate - sending preset for OSCO - HD 216399

-guidestar not found - doing a pointing check

UT 01:09 - on-target - WFS 1.15", Guider 1.3", DIMM 1.4"

UT 01:17 - aborting OSCO - re-rotator on w-unit causing problems

-LUCI rotator has to be > 0 degrees to avoid this issue.

-need to pick position angles to avoid this - from the Com Wiki:

ROTANGLE = POSANGLE - PARANGLE + 7.7

-so need to pick PAs and take into account the parallactic angle of target when observing it. Probably best to start with PA=180

UT 01:21- still waiting for re-rotator to home

UT 01:32 - still waiting for re-rotator to home

UT 01:59 - restarted w-unit

ESM

UT 02:34 - tried to send a preset with IDL as TTM but it went a ACE ACTIVE, called Doug, who sent the same thing and it worked - same account same OBS station, only difference is that I used TS=0 he did not use TS keyword

-configuring LUCI-1 for N375, imaging

First target AO 829, R=10 (used Xephem to find target, remember, its 'Xephem' not 'xephem'

UT 02:48 - still having issues with re-rotator (Doug on the line)

UT 02:50 - taking test expsure for R=10 star with [FeII]

-image is poorly collimated but near saturation, putting ND2 into place

ND2+[FeII] works well

UT 03:42 - still troubleshooting - G. Taylor on phone with Ale

Switching to backup SL

Going to binocular LUCI - will obtain M52 astrometric mask Rot ZPs with both LUCIs

-M52 first

LUCI-1 RotZP = -187.315

LUCI-2 RotZP = 176.85

UT 03:54 - pointing and collimation check

UT 4:00 - slewing to M52 LUCI-Bino

L1 0005 - field image

L1 0006 - target image

L1 0007 - thru mask

L2 0003 - field image

L2 0004 - target image

L2 0005 - thru mask

SIDENOTE: LUCI-2 RTD SHOWED THE WRONG LMS FILE BUT LUCI-1 RTD DID NOT (this is first use of RTD after mask exchange)

LUCI-2

using box 1 and box 6: X-shift = -2.44 pix, Y-shift =-1.45 pix, dtheta = -0.2997

Solution gives X-shift = 4.195pix, Y-shift = -2.651, dtheta=-0.361 -red errors (auto find)

Mis-identified #34 - picked a much brigher star NE of it., selected that and errors are 0.06 and 0.07 in RA/Dec

X-shift = 3.765 pix

Y-shift = -2.109 pix

dtheta = -0.374

LUCI-1

using box 1 and box 6: X-shift = 6.496 pix, Y-shift = 12.049 pix, dtheta = 0.062

Solution gives X-shift = 5.792 pix Yshift=-1.954 pix, dtheta = 0.07638 red errors (auto find)

Mis-identified #34 - picked a much brigher star NE of it. removed #38 it had 0.2 error

new solution:
X-shift = 5.339 pix

Y-shift = -1.422pix

dtheta = 0.069

UT 04:22 - taking confirmatory thru-slit images:

L1 0008 - stars look well centered in boxes

L2 0006 - stars look well centered in boxes

UT 04:24 - moving to TSIP Mask #1

UT 04:29 LUCI-2 MOS error (LUCI-1 in turnout = putting mask in FPU) elevation 51 rotator 188

Error while trying to move a mask from the focal plane unit to the turnout position at MOS unit "rmi://localhost:60002/T.

Focal plane open, picker arms open

Spindle is -10848 - which corresponds to Slot 15

Rotator is 341

Strain gauges are off, postconditions on FPU to Storage are OK. all masks are locked

Retainer is locked

Pre-conditions for translator are OK

Initializing the MOS unit

Init oK

Manually putting the mask into FPU

Its almost like the MOS forgotwhat it was doing halfway through....this was a twinned script, L1 completed just fine

UT 04:44 - resending LUCI-Bino TSIP Mask#1

UT 04:47 - same error! - its commanding the mask to be put away then throwing up an error - masks are locked

Putting in mask manually again, then will skip to next step

L1 0009 - field image

L2 0007 - field image

Manually putting mask in FPU, then skipping to line where it takes image

-Its doing it again, its forcing the mask in LUCI2 to be moved! inti MOS and putting mask in manually

will do LUCI-1 alignment in the meantime:

LUCI-1

Using box 1 and 6: dtheta = 0.0742, xshift = 6.022pix yshift=8.917pix

Fit is dtheta=0.154, xshift=0.85, yshift=-1.00, errors in yellow, checking stars:

Had to remove 31, 32, and 36. #31 looks directly above the star, whereas the others are all slightly shifted to the NE (propoer motion)

with those three removed:

dtheta = 0.427, xshift=-6.708pix yshift=-8.51pix

LUCI-2

Using box 1 and 6: dtheta = -0.2273, xshift=-11.219pix yshift=-0.7089pix

All stars gives dtheta=-0.0471 xhisft=-4.086pix yshift=-13.772 pix, yellor errors in RA

Removing Stars #31,32, and 36 gives:

dtheta = -0.051, xhisft=-4.338pix yshift=-13.537pix

UT 05:13 - hit GO in GUI to do offsets, but pause before image taken, manually takng exposure

Thru slits look good.

DONE - back to ESM

ESM Part Deux

UT 05:20 - authorizing LUCI-1 only

Pointing check

UT 05:32 Moving to AO 182 mH=9.34 - using [FeII] R=12.0

had wonkiness with sending IDL preset, this worked finally:

IDL> iif_register, /TCS
IDL> iif_preset, AO182, MODE=TTM, rs=0, gs=3,/left
IDL> iif_preset,/left

IDL> iif_runao

UT 05:37 - shell RIP (SX)

UT 05:42 - resending preset to AO 182

UT 05:43 - run ao

L1 0013 -test exposure 2.51sec [FeII] - peak 27k

2.51sec x 26 NDIT savemode=cube

UT 05:52 - L1 0014 - FWHM 3.18pix = 0.37"

UT 05:56 - L1 0015 - offsetting +5" +5" - FWHM 2.59pix = 0.30"

UT 05:59 - L1 0016 - offsetting -5" -5" - FWHM 2.99pix = 0.35"

UT 06:05 - moving to AO 247 - R=8 mH=6.25, GS=8, pa=180

DIMM 0.9"-1.1"

L1 0017 - test exposure 2.51 sec [FeII] - > 58k

L1 0018 - test exposure 2.51 sec [FeII] + ND2 - 3.2k

L1 0019 - -using 6 second DIT [FeII] + ND2 - 9k

UT 06:11 - L1 0020 - 6sec x 15 DIT - [FeII] + ND2 - FWHM 2.84pix = 0.33"

UT 06:13 - L1 0021 -5" -5" offset - DIMM 1.18" - FWHM 3.06pix = 0.036"

UT 06:16 - L1 0022 +5" +5" offset - DIMM 1.02"

UT 06:20 - moving to AO126, GS=5, PA=0 R=12.5, mH=8.67

L1 0023 - test exposure H-band 2.51 sec > 53k

L1 0024 - [FeII] 2.51 sec - 30k

L1 0025 - 2.51sec x 26 NDIT cube [FeII] - FWHM 3.91pix = 0.46"

L1 0026 - +5" +5" offset - FWHM 4.14pix = 0.49"

L1 0027 - -5" -5" offset - putting in ND2 - 2.51 sec test exposure

L1 0028 - -5" -5" offset ND2+[FeII] 6sec x 15 NDIT cub - FWHM 2.75pix = 0.32" DIMM 1.1"

L1 0029 - offsetting back to 0, 0 - 6x15 [FeII]+ND2 -

UT 06:43 - AO 187, GS=5, PA=0 R=15.9 mH=11.3

L1 0030 - test exposure [FeII]+ND2 2.51 sec ~100 counts

L1 0031 - 30DIT x 3 NDIT Cube [FeII]+ND2 - FWHM 2.32pix - 0.27"

L1 0032 - +10" +10" offset - 30DIT x 3NDIT Cube [FeII]+ND2 - 2.50pix = 0.29" DIMM =1.4"

-Note: I thought I had put in H-band filter, but had left the [FeII]+ND2 in pace - still got enough counts to make a decent measurement

UT 06:59 - moving to AO 259, GS=2, PA=0 R=7.8 mH=5.61

L1 0033 - test exposure [FeII]+ND2 2.51sec -11k

L1 0034 - [FeII]+ND2 - 2.51x26 cube FWHM 3.03pix = 0.36"

L1 0035 - [FeII]+ND2 - 2.51x26 cube (didn't offset) FWHM 3.25pix = 0.38" (didn't offset)

L1 0036 - +10" +10" offset - [FeII]+ND2 - 2.51x26 cube FWHM 2.80 pix = 0.33"

L1 0037 - offsetting -10" -10" [FeII]+ND2 - 2.51x26 cube FWHM 3.20pix = 0.38" DIMM 1.2"

Completed single star tests

UT 07:18 - authorizing LUCI-Bino

Will go to M34 - R=12.4 AO Ref star - SX-ESM, DX-SL

UT 07:25 - manually putting mask in LUCI-2 away, going to No Mask in Use

UT 07:30 - sending bino M34 script

UT 07:39 - loops closed - integrating

L1 0038-0067

L2 0011-0040

UT 07:41 - copointing limit error due to large offsets to sky

- need to disable them - IT 6553

-Continuing scripts

FWHM first image 3.3-3.8 pix --> 0.39"-0.45" DIMM 1.5"

UT 07:59 - gettng scattered light on clouds off the Moon, might be affecting WFS

UT 08:03 - switched to H-band, improvement went from 200% to 30% relative to SL on DX

DX guidecamera suggests ~1 mag of extinction

UT 08:33 Finished M34 (El 44)

UT 08:33 - Moving to Arp 143 - AO Ref star missed CCD47

UT 08:37 - doing a pointing check

UT 08:48 resending Arp 143 Bino preset (also had to fix script)

FWHM ~0.55" ESM vs 0.8" SL

L1 0068-0079

L2 0041-0052

L1 0071 - outside of Patrol Field do not use

-redid script with different offsets starting from position 3

New Script PA=180 - this gets SF blobs on top of array

UT 09:26 - may have reached a limiting point, background is too high to close AO loop

-failed on 0,0 offset

Poor Weather H-band Library

UT 09:30 - Configuring for LUCI-1 only to avoid over-using AGW-4 on LUCI-2

-putting LUCI-2 into safe mode

Telluric 49Aur (for HD 39225 and HD 42475) first guidestar R=13 will try

4 mags of extinction

UT 09:47 - starting spectroscopy

L1 0080-0090

UT 10:01 - moving to HD 39225 M1III

UT 10:07 starting spectroscopic data

L1 0091-0093 - Acquisition

-lost guidestar - could be because it was at edge of patrol field, picking something closer in and as bright

UT 10:16 - starting spectroscopy - object might be too bright

L1 0094-0104

UT 10:30 - moving to HD 42475 M2I

UT 10:37 - starting spectroscopy

L1 0105-0116 - may be bad, clouds highly variable - but may not have been thin enough

UT 10:57 - moving to Telluric 23LMi (for HD 83787)

UT 11:05 - starting spectroscopy

L1 0117- 0124 - competed telluric

UT 11:14 - clouds may be thinning enough to try mergers

Backup Program - CO Bandhead Mergers

UT 11:15 - going to HD99966- Telluric for NGC 3921

UT 11:22 - starting spectra

L1 0125-0130

-definitely thinned out overhead, not enough extinction for H-band library anymore

UT 11:29- Moving to NGC 3921

Guider shows 0.63"

UT 11:39 - starting spectra

L1 0131-0142

UT 13:05 - finished - closing up to take calibratons

G210 K CWL 2.27 0.5" slit L1 Cals

L1 0143-0147 - flats lights off

L1 0148-0152 - flats lights on

L1 0153-0154 - arcs lights off

L1 0155-0156 - neon lights on

L1 0157-0158 - argon lights on

L1 0159-0160 - xenon lights on

G210 H CWL 1.65 0.5" slit L1 Cals

L1 0161-0165 - flats lights off

L1 0166-0170 - flats lights on

L1 0171-0172 - lights off argon

L1 0173-0174 - lights on argon

L1 0175-0176 - lights off xenon

L1 0177-0178 - lights on xenon

LUCI-1 and LUCI-2 N375 Imaging Flats, K, H, and J

L1 0179-0183 L2 0053-0057 J lights off

L1 0184-0188 L2 0058-0062 J lights on

L1 0189-0193 L2 0063-0067 H lights off

L1 0194-0198 L2 0068-0072 H lights on

L1 0199-0202 L2 0073-0077 K lights off

L1 0203-0207 L2 0077-0084 K lights on

UT 14:10 - Cals complete - LUCIs in safe config

-- BarryRothberg - 10 Dec 2019
Topic revision: r2 - 11 Dec 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback